OpenStudy (anonymous):

rationalize the denominator and simplify (2√7+3)/(√7+2)

OpenStudy (anonymous):

When you want to rationalize the denominator you want to multiply it by the numerator and denominator by the conjugate \[\sqrt{7}-2\]

OpenStudy (anonymous):

\[\frac{ (2\sqrt{7}+3 )}{ (\sqrt{7}+2 )} \times \frac{ \sqrt{7}-2 }{ \sqrt{7}-2 }\]

OpenStudy (anonymous):

why is it negative? shouldnt it be positive 2?

OpenStudy (anonymous):

oh opps just kidding the bottom cancels out

OpenStudy (anonymous):

To rationalize we use the conjugate which simply means we change the sign so notice how we want to rationalize the denominator so we multiply the top and bottom by sqrt(7)-2 we change the + to - this way we can get an expression as such \[a^2-b^2\]

OpenStudy (anonymous):

Yeah!

OpenStudy (usukidoll):

"When you want to rationalize the denominator you want to multiply it by the numerator and denominator by the conjugate " signs need to be switched.. since you have a + on the denominator we have to multiply by sqrt{7}-2

OpenStudy (anonymous):

ahaha I understand now! thank you both!